Velocity of 54m/s acceleration uniformly at rate of 2ms velocity reaches 72m/s




Calculate the distance travelled from the place where the acceleration began




The time taken to cover this distance

1
Expert's answer
2021-11-16T09:51:41-0500

Since the distance traveled is "S =\frac {V _ 1 ^ 2-V _ 0 ^ 2} {2\times a}" , where "V_1" is the final speed, "V_0" is the initial, "a" is the acceleration, then "S =\frac {72 ^ 2-54 ^ 2} {2\times 2} = 567" m;

Since by another formula the distance is "S = V _ 0\times t +\frac {a\times t ^ 2} {2}" , we will express from this equation "t" :

"\frac {a\times t^2}{2}+V_0\times t - S=0" ,

"\frac {2\times t^2}{2}+54\times t - 567=0" ,

"t=9" s.